In an isoceles triangle, the length of the third side is 25% of one of the equal sides. Find the length of each side when perimeter is equal to 135 ft.

Answer by TimothyLamb(4379)

x = length of third side
y = length of each equal side
---
x + 2y = 135
x = 0.25y
---
put the system of linear equations into standard form
---
x + 2y = 135
x - 0.25y = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = length of third side = 15 ft
y = length of each equal side = 60 ft
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---